# Zero-shot Generalization in Inventory Management: Train, then Estimate and Decide

### v2

@@ Appendix D Estimate - Decide phases for periodic review inventory control
  • At the start of a period, we place orders using only information observed up to that time; the decision does not depend on the yet-unrealized demand, and thus the available-stock (censoring) threshold is fixed before demand is realized (cf. Huh et al.,, 2011).
@@
When we start collecting observations, related unknown parameters are estimated. For the unknown demand distribution, we utilize the Kaplan-Meier estimator (Kaplan and Meier,, 1958). While initially proposed for estimating the survival function from lifetime data, the Kaplan-Meier estimator is also adopted in inventory problems. Interested readers may refer to Huh et al., (2011); Lyu et al., (2024) for the usage of it to build an empirical cumulative distribution function (CDF) from the censored demand observations. While we depart from them in having a cyclic demand distribution, the procedure stays the same as we know the cycle length $K$, see Assumption 3. So we repeat the process for each subset of demands within a cycle. We then derive the estimated mean demand $\hat{\mu}$ and the standard deviation of the demand $\hat{\sigma}$ from the empirical CDF to create and update the demand distribution estimate $\hat{\boldsymbol{\zeta}}=(\hat{\mu_{0}},\hat{\sigma_{0}},\dots,\hat{\mu}_{K-1},\hat{\sigma}_{K-1})$.
@@
Fix $j$. Within cycle $j$ the demands are i.i.d. (discrete) and, by Assumption 3, the censoring level $OH_{t}+q_{t}+q_{t}^{0}$ is chosen from past information, so censoring is independent of the demand. Therefore, the conditions of Theorem 2 in Huh et al., (2011) hold for cycle $j$, which yields $\sup_{x\in\{0,\dots,D_{\max}\}}|\widehat{F}_{t,j}(x)-F_{j}(x)|\to 0$ almost surely. ∎
@@ Appendix E Experimental Setup
We normalize the holding cost to $h=1.0$, and set $p_{\min}=2.0$ since in lost sales systems, penalty costs are typically higher than holding costs; otherwise, holding no inventory would be favorable (cf. Zipkin,, 2008; Xin,, 2021). We cap the penalty cost at $p_{\max}=100.0$ and lead time at $L_{max}=10$, as in regimes with very high penalty costs and long lead times, the capped base-stock policy is known to be asymptotically optimal (Xin,, 2021), reducing the benefit of a sophisticated dynamic policy. For the demand distribution, we use $\mu_{\min}=2.0$ to avoid sparse demand levels, and set $\mu_{\max}=12.0$ for computational tractability, since the size of the action space grows with demand, and larger values would require extensive sampling (§5.1) and significantly higher compute. A possible extension would be to introduce a low/high demand parameter (such as $l$ which shows whether orders cross or not), where in high-demand settings, a single action corresponds to a batch order with predefined quantities. We set $K_{\max}=7$ to reflect weekly demand cycles.
Table 6 presents the set of hyperparameters used to train the Super-DCL algorithm in our numerical experiments. For a comprehensive overview of the hyperparameters utilized, we direct readers to Temizöz et al., (2025). We employed the identical neural network architecture described in their study and adhered to the selection criteria for sampling and simulation-related hyperparameters outlined therein. To this end, we increased the number of samples $N$ and the number of approximate policy iterations $n$ for performance gains, and decreased the depth of the simulations $H$ and simulation budget per state-action pairs $M$ for computational efficiency. We refrained from further hyperparameter tuning. Departing from that study, we introduce two new hyperparameters: $R$, the number of samples collected for a parameterization; $P$, the maximum number of promising actions for simulations (see Danihelka et al.,, 2022, for an example usage).
@@ Appendix G Instances of the Numerical Experiments
We then fit a demand distribution by using the two moments as described by Adan et al., (1995).
@@
For Case 3, we consider the following demand distributions: for $K=1$, $\mu_{1}=5.0$; for $K=3$, $\mu_{1}=8.0,\mu_{2}=10.0,\mu_{3}=6.0$; for $K=5$, $\mu_{1}=11.0,\mu_{2}=3.0,\mu_{3}=5.0,\mu_{4}=8.0,\mu_{5}=6.0$; for $K=7$, $\mu_{1}=3.0,\mu_{2}=5.0,\mu_{3}=5.0,\mu_{4}=7.0,\mu_{5}=7.0,\mu_{6}=10.0,\mu_{7}=10.0$. The corresponding standard deviations are chosen according to the following distributions: $\sigma_{1}-Poisson,\sigma_{2}-Geometric,\sigma_{3}-NegativeBinomial,\sigma_{4}-Binomial,\sigma_{5}-Poisson,\sigma_{6}-Geometric,\sigma_{7}-NegativeBinomial$. We then fit a demand distribution by using two moments as described by Adan et al., (1995).
@@ Appendix H Per-Instance Training and Instance Selection
We quantify the performance and operational trade-off when facing a choice between training and deploying a GCA with ZSG ability and training individual agents tailored to fixed, known parameters. We therefore collect 35 instances spanning all three experimental cases and train individual DRL agents for each instance using the same neural network architecture and training pipeline as GC-LSN (identical hyperparameters; only the number of training samples is adjusted to ensure per-instance convergence; same hardware).
For training individual DRL agents tailored to specific parameterizations, we employ the DCL algorithm (Temizöz et al.,, 2025). We use the same hyperparameters reported in Table 6, except for the number of samples, which is set to $N=100000$. Note that the hyperparameters $R$ and $P$ are specific to the Super-DCL algorithm, which applies to training Super-MDPs, and are therefore not considered here. Total training time for all agents was $\sim$ 6 hours.
@@
Case 1: We consider 16 problem instances using a full factorial design with the following parameter ranges: mean demand $\mu\in\{5.0,10.0\}$, penalty cost $p\in\{9.0,69.0\}$, deterministic lead time $\in\{4,8\}$. For each mean demand value, the standard deviation is selected such that the demand distribution follows either Poisson or geometric.
@@
We benchmark GC-LSN against these individual agents under known parameterizations. In summary, GC-LSN incurs a slightly higher cost on average, with a mean relative cost gap of only $0.2\%$ versus per-instance agents under known parameterizations. Given this small gap and the retraining burden of per-instance models as tasks scale or parameters change, and the fact that per-instance training presumes known parameters (hence is infeasible when they are unknown), GC-LSN’s ZSG capability offers a practical advantage in settings with many tasks or frequent updates.
@@ Appendix I Restoring Lipschitz Property in Super-MDPs
Fix $(\mathbf{s},a,\theta)$ and choose $\beta^{\prime}=\mathrm{FR}(\mathbf{s},a,(\theta,\beta^{\prime}))$ (e.g., by taking a state/action where the fill rate equals the threshold). Then for any $\varepsilon>0$,
@@
If $\beta$ is truly continuous, however, then the probable parameter space $\hat{\mathcal{P}}$ can no longer be a strict superset of the true space $\mathcal{P}$: exact coverage would require infinitely many grid points. In practice, achieving adequate coverage of a continuous $\beta$ requires choosing a sufficiently fine discretization level. Larger $Z$ improves the approximation and reduces the Lipschitz constant’s looseness but also increases the dimensionality of the training space and hence the sampling budget needed to cover it uniformly. Thus, while discretization restores the Lipschitz property, it introduces a trade-off between fidelity to the true continuous space and computational feasibility.

### v3

@@ Appendix D Estimate - Decide phases for periodic review inventory control
  • At the start of a period, we place orders using only information observed up to that time; the decision does not depend on the yet-unrealized demand, and thus the available-stock (censoring) threshold is fixed before demand is realized (cf. Huh et al., 2011).
@@
When we start collecting observations, related unknown parameters are estimated. For the unknown demand distribution, we utilize the Kaplan-Meier estimator (Kaplan and Meier, 1958). While initially proposed for estimating the survival function from lifetime data, the Kaplan-Meier estimator is also adopted in inventory problems. Interested readers may refer to Huh et al. (2011); Lyu et al. (2024) for the usage of it to build an empirical cumulative distribution function (CDF) from the censored demand observations. While we depart from them in having a cyclic demand distribution, the procedure stays the same as we know the cycle length $K$, see Assumption 3. So we repeat the process for each subset of demands within a cycle. We then derive the estimated mean demand $\hat{\mu}$ and the standard deviation of the demand $\hat{\sigma}$ from the empirical CDF to create and update the demand distribution estimate $\hat{\boldsymbol{\zeta}}=(\hat{\mu_{0}},\hat{\sigma_{0}},\dots,\hat{\mu}_{K-1},\hat{\sigma}_{K-1})$.
@@
Fix $j$. Within cycle $j$ the demands are i.i.d. (discrete) and, by Assumption 3, the censoring level $OH_{t}+q_{t}+q_{t}^{0}$ is chosen from past information, so censoring is independent of the demand. Therefore, the conditions of Theorem 2 in Huh et al. (2011) hold for cycle $j$, which yields $\sup_{x\in\{0,\dots,D_{\max}\}}|\widehat{F}_{t,j}(x)-F_{j}(x)|\to 0$ almost surely. ∎
@@ Appendix E Experimental Setup
We normalize the holding cost to $h=1.0$, and set $p_{\min}=2.0$ since in lost sales systems, penalty costs are typically higher than holding costs; otherwise, holding no inventory would be favorable (cf. Zipkin, 2008; Xin, 2021). We cap the penalty cost at $p_{\max}=100.0$ and lead time at $L_{max}=10$, as in regimes with very high penalty costs and long lead times, the capped base-stock policy is known to be asymptotically optimal (Xin, 2021), reducing the benefit of a sophisticated dynamic policy. For the demand distribution, we use $\mu_{\min}=2.0$ to avoid sparse demand levels, and set $\mu_{\max}=12.0$ for computational tractability, since the size of the action space grows with demand, and larger values would require extensive sampling (§5.1) and significantly higher compute. A possible extension would be to introduce a low/high demand parameter (such as $l$ which shows whether orders cross or not), where in high-demand settings, a single action corresponds to a batch order with predefined quantities. We set $K_{\max}=7$ to reflect weekly demand cycles.
Table 6 presents the set of hyperparameters used to train the Super-DCL algorithm in our numerical experiments. For a comprehensive overview of the hyperparameters utilized, we direct readers to Temizöz et al. (2025). We employed the identical neural network architecture described in their study and adhered to the selection criteria for sampling and simulation-related hyperparameters outlined therein. To this end, we increased the number of samples $N$ and the number of approximate policy iterations $n$ for performance gains, and decreased the depth of the simulations $H$ and simulation budget per state-action pairs $M$ for computational efficiency. We refrained from further hyperparameter tuning. Departing from that study, we introduce two new hyperparameters: $R$, the number of samples collected for a parameterization; $P$, the maximum number of promising actions for simulations (see Danihelka et al., 2022, for an example usage).
@@ Appendix G Instances of the Numerical Experiments
We then fit a demand distribution by using the two moments as described by Adan et al. (1995).
@@
For Case 3, we consider the following demand distributions: for $K=1$, $\mu_{1}=5.0$; for $K=3$, $\mu_{1}=8.0,\mu_{2}=10.0,\mu_{3}=6.0$; for $K=5$, $\mu_{1}=11.0,\mu_{2}=3.0,\mu_{3}=5.0,\mu_{4}=8.0,\mu_{5}=6.0$; for $K=7$, $\mu_{1}=3.0,\mu_{2}=5.0,\mu_{3}=5.0,\mu_{4}=7.0,\mu_{5}=7.0,\mu_{6}=10.0,\mu_{7}=10.0$. The corresponding standard deviations are chosen according to the following distributions: $\sigma_{1}-Poisson,\sigma_{2}-Geometric,\sigma_{3}-NegativeBinomial,\sigma_{4}-Binomial,\sigma_{5}-Poisson,\sigma_{6}-Geometric,\sigma_{7}-NegativeBinomial$. We then fit a demand distribution by using two moments as described by Adan et al. (1995).
@@ Appendix H Per-Instance Training and Instance Selection
We quantify the performance and operational trade-off when facing a choice between training and deploying a GCA with ZSG ability and training individual agents tailored to fixed, known parameters. We therefore collect 35 instances spanning all three experimental cases and train individual DRL agents for each instance using the same neural network architecture and training pipeline as GC-LSN (with identical hyperparameters; only the number of training samples is adjusted to ensure per-instance convergence; the same hardware).
For training individual DRL agents tailored to specific parameterizations, we employ the DCL algorithm (Temizöz et al., 2025). We use the same hyperparameters reported in Table 6, except for the number of samples, which is set to $N=100000$. Note that the hyperparameters $R$ and $P$ are specific to the Super-DCL algorithm, which applies to training Super-MDPs and are therefore not considered here. Total training time for all agents was $\sim$ 6 hours, with each individual agent requiring less than 15 minutes to train.
@@
Case 1: We consider 16 problem instances using a full factorial design with the following parameter ranges: mean demand $\mu\in\{5.0,10.0\}$, penalty cost $p\in\{9.0,69.0\}$, and deterministic lead time $\in\{4,8\}$. For each mean demand value, the standard deviation is selected so that the demand distribution follows either a Poisson or geometric distribution.
@@
We benchmark GC-LSN against these individual agents under known parameterizations. In summary, GC-LSN incurs a slightly higher cost on average, with a mean relative cost gap of only $0.2\%$ versus per-instance agents under known parameterizations. Given this small gap, the retraining burden of per-instance models as tasks scale or parameters change, and the fact that per-instance training presumes known parameters (hence is infeasible when they are unknown), GC-LSN’s ZSG capability offers a practical advantage in settings with many tasks or frequent updates.
@@ Appendix I Restoring Lipschitz Property in Super-MDPs
Fix $(\mathbf{s},a,\theta)$ and choose $\beta^{\prime}=\mathrm{FR}(\mathbf{s},a,(\theta,\beta^{\prime}))$ (e.g., by taking a state/action where the fill rate equals the threshold). Then, for any $\varepsilon>0$,
@@
If $\beta$ is truly continuous, however, then the probable parameter space $\hat{\mathcal{P}}$ can no longer be a strict superset of the true space $\mathcal{P}$: exact coverage would require infinitely many grid points. In practice, achieving adequate coverage of a continuous $\beta$ requires choosing a sufficiently fine discretization level. Larger $Z$ improves the approximation and reduces the looseness of the Lipschitz constant, but it also increases the dimensionality of the training space and hence the sampling budget needed to cover it uniformly. Thus, while discretization restores the Lipschitz property, it introduces a trade-off between fidelity to the true continuous space and computational feasibility.